# Run-Time Monitoring of ERTMS/ETCS Control Flow by Process Mining

### v1

@@ Abstract
Ensuring the resilience of computer-based railways is increasingly crucial to account for uncertainties and changes due to the growing complexity and criticality of those systems. Although their software relies on strict verification and validation processes following well-established best-practices and certification standards, anomalies can still occur at run-time due to residual faults, system and environmental modifications that were unknown at design-time, or other emergent cyber-threat scenarios. This paper explores run-time control-flow anomaly detection using process mining to enhance the resilience of ERTMS/ETCS L2 (European Rail Traffic Management System / European Train Control System Level 2). Process mining allows learning the actual control flow of the system from its execution traces, thus enabling run-time monitoring through online conformance checking. In addition, anomaly localization is performed through unsupervised machine learning to link relevant deviations to critical system components. We test our approach on a reference ERTMS/ETCS L2 scenario, namely the RBC/RBC Handover, to show its capability to detect and localize anomalies with high accuracy, efficiency, and explainability.
Keywords:
train control systems dependability fault-tolerance software reliability radio block center.
@@ 1 Introduction
The European Rail Traffic Management System (ERTMS) [^1] https://www.era.europa.eu/domains/infrastructure/european-rail-traffic-management-system-ertms_en is a European standard setting the specification of both architecture and functions for interoperable, efficient, and dependable railways. The European Train Control System (ETCS) standardizes the Automatic Train Protection (ATP) subsystem [16]. ERTMS/ETCS has different levels of operation, which determine how the on-board equipment of trains exchanges key information with the trackside subsystems. In this paper, we address ERTMS/ETCS L2, which uses wireless communication between on-board equipment and the Radio Block Center (RBC), based on the Euroradio protocol, to ensure continuous signalling. Train position is determined through Eurobalises and sent to the RBC via the Position Reports (PR); RBC computes and transmits to the trains the so-called Movement Authority (MA), when full supervision is active (i.e., after start of mission and in case of no failures).
@@
ERTMS/ETCS L2 must undergo strict verification and validation processes according to well-established best-practices, internal standards, and certification requirements. Those include extensive requirements engineering, modeling and formal verification also using model-checking [3, 9, 14, 16, 20, 7], as well as simulation and testing [31, 32, 2, 15, 24, 30]. However, the coverage of those activities can never be complete due to several factors, such as: (1) the usage of natural language specification that might have ambiguities; (2) the high level of complexity undermining full formal verification and testing coverage, (3) heterogeneity in components developed by different manufacturers using diverse software versions and implementations. In such a scenario, residual or interaction faults, unexpected or uncontrolled modifications, system or environmental uncertainties, different specification interpretation or implementation across multiple manufacturers, emergent cyber-threats, as well as any other “unknown unknowns”, pose threats that might affect correct execution of ERTMS/ETCS L2 procedures and generate failures. Generally speaking, functional, environmental and technological changes during system operation require appropriate means to ensure resilience, i.e., the persistence of dependability when facing changes [21].
@@
To address the problem of ensuring dependability in the presence of unexpected events possibly corrupting ERTMS/ETCS L2 control flow, we propose an approach for run-time monitoring and anomaly detection through process mining, which encompasses a set of explainable algorithms to model the so-called normative behavior (i.e., process discovery), and to check any deviations during system operation (i.e., conformance checking) [1]. In fact, process mining allows:
@@
  • Monitoring the execution of ERTMS/ETCS L2 procedures under normal conditions and modeling their behavior through process discovery;
@@
  • Collecting run-time control-flow of ERTMS/ETCS L2 and checking execution correctness against the normative model through conformance checking.
Such an approach enables resilience to changes and uncertainties as it combines data-driven insights with model-based analyses, taking concrete steps toward the development of an explainable run-time monitor that learns from actual system operation.
Process mining has been applied for anomaly detection in several domains, including business processes of organizations [22, 35], software applications [11, 25], computer networks [28, 18], and cyber-physical systems [23, 34]. Notably, process mining has been also applied using an ERTMS/ETCS dataset generated from a high-level description of one use-case scenario in [12, 35]. However, previous works did not focus on a methodology for improving the resilience of ERTMS/ETCS through run-time monitoring and localization of control-flow anomalies using process mining, which is the scope of this paper.
@@ 2.1 Preliminary Definitions
Traditional definitions of dependability typically emphasize either “the ability to deliver service that can justifiably be trusted” or “the ability to avoid service failures that are more frequent and more severe than is acceptable” [4]. These definitions encompass a broad range of strategies for achieving dependability, including fault prevention, fault tolerance, fault removal, and fault forecasting. However, they do not explicitly address a key concept that is central to this paper: resilience. Resilience refers to a system’s ability to continue delivering correct service in the face of changes in its run-time environment, including functional, environmental, and technological changes [21].
To achieve resilience in modern computer-based railways, including those based on ERTMS/ETCS L2, we propose to add specific software monitors performing run-time control-flow anomaly detection. Anomaly detection is a broad research area that aims to find “patterns in data that do not conform to a well defined notion of normal behavior" [8]. Different types of anomaly detection address different types of anomalies. In this paper, we address control-flow anomalies, namely those anomalies that manifest as unknown, skipped, and wrongly-ordered activities of event logs from computer-based systems [35]. In addition, we aim to perform control-flow anomaly detection at run-time, including when the system is operating in non perfectly known/predictable conditions, hence subject to functional, environmental, and technological uncertainties. Finally, it is worth mentioning that we propose an unsupervised approach to anomaly detection and localization, as we aim to characterize normal behavior from reference execution traces of the target system, and localize anomalies based on the model-based diagnoses we perform against a reference process model of the software.
@@ 2.2 ERTMS/ETCS L2
ERTMS/ETCS is characterized by a certain level of operation, from 1 to 3, which defines the degree of automation performance, and the required trackside and on-board equipment [16].
@@
In this paper, we focus on the most successful level of operation to date, which is level 2. At this level, the on-board system collects the train’s position through Eurobalises, which are electronic beacons installed across the tracks. The position report is forwarded to the RBC, which is responsible for train separation; the RBC elaborates the MA based on track occupation. Fig. 1 shows a common use-case in ERTMS/ETCS, namely the MA request and provision to supervise train traffic. This use-case can be implemented within the Start-of-Mission procedure, which is one of the procedures documented in specification SUBSET-026; such a subset prescribes the behavior of ERTMS/ETCS constituents in all reference operational scenarios. Each step, labeled 1 to 7, requires either exchanging information between the on-board equipment of the requesting and preceding trains with the RBC, or using the on-board modules to perform some kind of data processing. The most important task performed on-board is the calculation of the so-called braking curve (also known as dynamic speed profile).
We assume that several types of run-time anomalies may occur across the many modules that compose ERTMS/ETCS L2, such as unexpected train data exchanged between the Driver Machine Interface (DMI) and the RBC through the Radio Transmission Module (RTM), mispositioning of balises, and issues in establishing connections with the RBC [31, 32, 6]. These anomalies may propagate across different components, and may influence other key modules, such as the European Vital Computer (EVC), which is the on-board component responsible for performing critical operations, such as calculating the braking curve.
@@ 2.3 Process Mining and Machine Learning
Process mining “aims to improve operational processes through the systematic use of event data” by mainly two types of algorithms: process discovery and conformance checking [1]. Process discovery attempts to build a behavioral model from historical event data collected by monitoring a reference application or process. Conformance checking aligns new event data to the behavioral model to diagnose any deviations from the control flow captured by the model. In addition, conformance checking can be performed online, i.e., it can be applied to streaming event data through, e.g., window-based approaches and prefix alignments [5, 37].
Machine learning allows encoding data patterns into models. Traditional machine learning approaches capture patterns in data through rule-based learning, probabilistic modeling, linear and non-linear modeling, and clustering [36]. Advanced techniques include deep learning, which implements (deep) artificial neural networks mimicking the functioning of the human brain [13]. Machine learning algorithms also differ based on their learning paradigm (supervised, unsupervised, and semi-supervised) and degree of explainability [26].
Previous works have shown that conformance checking can be combined with unsupervised machine learning for explainable control-flow anomaly detection. The approach proposed in this paper is inspired by the work reported in reference [35], and proposes performing unsupervised run-time monitoring of ERTMS/ETCS L2 control flow through the combination of conformance checking and clustering.
@@ 3.1.1 Offline ERTMS/ETCS Execution
The first step begins with an offline ERTMS/ETCS execution, where a test suite is applied to a simulation environment or an actual system implementation. In the case of simulating the system, several types of high-level events can be collected. For example, di Tommaso et al. [31] simulated the Track Ahead Free (TAF) procedure with a trackside simulation environment. In this procedure, the RBC sends a TAF request to the train driver through the DMI to possibly grant a full supervision MA. Gaspari et al. [15] exemplified another remarkable simulation approach, involving the execution of abstract state machines representing several ERTMS/ETCS L3 scenarios to test the correctness of their models. The execution traces are collected to model the software behavior. We define a set of $k$ traces as $\Sigma=\{\sigma_{1},\sigma_{2},\dots,\sigma_{k}\}$, where each trace $\sigma$ is an ordered sequence of events: $\sigma=\langle e_{1},e_{2},\dots,e_{|\sigma|}\rangle\in\Sigma$, with $e_{j}$ representing the $j$-th event. The definition of an event depends on the level of abstraction used in simulating ERTMS/ETCS L2. Since these systems involve distributed software components communicating over a railway network, events may correspond to specific software procedures executed by those components. This is a realistic representation, as safety-critical ETCS software is usually implemented in C/C++ [3, 9, 20].
@@ 3.1.2 Model Building
After the collection of execution traces, the model building activity applies a process discovery algorithm $\gamma$ to find an execution model $M$ from the set of traces representing the actual behavior of the system, i.e., $\gamma(\Sigma)=M$. There are many process discovery algorithms classified by their different learning approaches, including footprint-based and directly-follows-graph-based approaches that uncover local relationships between events (e.g., the $\alpha$-miner and split miner), divide-and-conquer approaches that iteratively find relationships starting from the most general one (e.g., the inductive miner and partially ordered workflow language miner), and region-based approaches that obtain so-called regions from either state-based or language-based represesentations of the input log (e.g., the integer linear programming-based miner) [1]. Each process discovery algorithm has its own representational bias, which influences the quality of the resulting execution model. The execution model can be, e.g., a Petri net, a business process modeling notation model, or a process tree. The target formalism depends on the requirements of the applications. For example, a Petri net may be better suited for safety-critical environments due to its formal semantics. Hence, we consider the Petri net as the target formalism in the following.
@@ 3.2.1 Online ERTMS/ETCS Execution
The second step is initiated by the online execution of ERTMS/ETCS with an unknown input. The real process is monitored throughout its execution. Such run-time monitoring can be performed in different ways. A plausible mechanism would be source-code instrumentation by well-defined logging rules, as proposed in [10]. This technique places logging rules to record specific events occurring in a program, such as interactions between objects, procedure calls, and the invocation of microservices. Its feasibility is due to the availability of source code in ERTMS/ETCS L2 implementations, as these systems require thorough white-box analyses to certify their safety levels. Fig. 3 shows an example deployment diagram where the on-board and trackside subsystems communicate with each other and send run-time events to a monitor server. These run-time events configure a partial trace, and are checked against the execution model by a fault detector using a conformance checker that implements process mining techniques.

### v2

@@ Abstract
Ensuring the resilience of computer-based railways is increasingly crucial to account for uncertainties and changes due to the growing complexity and criticality of these systems. Although their software relies on strict verification and validation processes following well-established best-practices and certification standards, anomalies can still occur at run-time due to residual faults, system and environmental modifications that were unknown at design-time, or other emergent cyber-threat scenarios. This paper explores run-time control-flow anomaly detection using process mining to enhance the resilience of ERTMS/ETCS L2 (European Rail Traffic Management System / European Train Control System Level 2). Process mining allows learning the actual control flow of the system from its execution traces, thus enabling run-time monitoring through online conformance checking. In addition, anomaly localization is performed through unsupervised machine learning to link relevant deviations to critical system components. We test our approach on a reference ERTMS/ETCS L2 scenario, namely the RBC/RBC Handover, to show its capability to detect and localize anomalies with high accuracy, efficiency, and explainability.
@@ 1 Introduction
The European Rail Traffic Management System (ERTMS) [^1] https://www.era.europa.eu/domains/infrastructure/european-rail-traffic-management-system-ertms_en is a European standard setting the specification of both architecture and functions for interoperable, efficient, and dependable railways. The European Train Control System (ETCS) standardizes the Automatic Train Protection (ATP) subsystem [ghazel2014formalizingertmsetcsspecs]. ERTMS/ETCS has different levels of operation, which determine how the on-board equipment of trains exchanges key information with the trackside subsystems. In this paper, we address ERTMS/ETCS L2, which uses wireless communication between on-board equipment and the Radio Block Center (RBC), based on the Euroradio protocol, to ensure continuous signalling. Train position is determined through Eurobalises and sent to the RBC via the Position Reports (PRs); RBC computes and transmits to the trains the so-called Movement Authority (MA), when full supervision is active (i.e., after start of mission and in case of no failures).
@@
ERTMS/ETCS L2 must undergo strict verification and validation processes according to well-established best-practices, internal standards, and certification requirements. These include extensive requirements engineering, modeling and formal verification also using model-checking [angeletti2010safetycriticalsoftwarecoverageanalysis, cimatti2012ertmsl2verificationvalidation, flammini2014multiformalismapproach, ghazel2014formalizingertmsetcsspecs, karg2016modeldrivenseopenetcs, campanile2022textbasedmetrics], as well as simulation and testing [ditommaso2005anomaliessimulation, valdivia2017etcssafetytesting, ameurbolifa2019verifyingetcssoftware, gaspari2019hybrideuropeanrailtrafficmgms, nardone2020systemlevelfunctionaltesting, su2022safetyinterlocking]. However, the coverage of these activities can never be complete due to several factors, such as: (1) the usage of natural language specification that might have ambiguities; (2) the high level of complexity undermining full formal verification and testing coverage, (3) heterogeneity in components developed by different manufacturers using diverse software versions and implementations. In such a scenario, residual or interaction faults, unexpected or uncontrolled modifications, system or environmental uncertainties, different specification interpretation or implementation across multiple manufacturers, emergent cyber-threats, as well as any other “unknown unknowns”, pose threats that might affect the correct execution of ERTMS/ETCS L2 procedures and generate failures. Generally speaking, functional, environmental and technological changes during system operation require appropriate means to ensure resilience, i.e., the persistence of dependability when facing changes [laprie2005resilience].
@@
To address the problem of ensuring dependability in the presence of unexpected events possibly corrupting ERTMS/ETCS L2 control flow, we propose an approach for run-time monitoring and anomaly detection through process mining, which encompasses a set of explainable algorithms to model the so-called normative behavior (i.e., process discovery), and to check any deviations during system operation (i.e., conformance checking) [aalst2022pmhb]. In fact, process mining allows:
@@
  • Modeling the execution of ERTMS/ETCS L2 procedures under normal conditions through process discovery;
@@
  • Collecting run-time ERTMS/ETCS L2 control-flow and checking execution correctness against the normative execution model through conformance checking.
The characterization of ERTMS/ETCS L2 procedures under normal conditions from actual execution traces helps bridge the gap between specifications/models and actual system implementations. Although model-driven engineering is a widely adopted practice when developing safety-critical systems, ensuring strict consistency between system models and their software implementations is challenging due to the system’s complexity and the concurrent operation of multiple developers [weidmann2018tolerant, trols2019multifaceted]. As a result, some degree of divergence between specifications/models and the actual software is often unavoidable. This makes the characterization of the system’s control flow from execution traces particularly valuable. In addition, while traditional rule-based methods (e.g., decision trees) provide a static representation of if–else rules [witten2016mltools] and more complex machine learning methods (e.g., recurrent neural networks) either lack an explicit representation of control-flow patterns or the ability to provide a process-based, trustworthy explanation of the deviations from prescriptive behavior [rawal2022trustworthyai, aalst2022pmhb], process mining explicitly allows capturing behavioral patterns — such as concurrent and exclusive control flow — through process discovery algorithms. Thereby, process discovery allows building prescriptive process models using well-known and interpretable formalisms, such as the Petri net. Finally, the approach enables resilience to changes, uncertainties, and any “unknown unknowns” as it combines data-driven insights with model-based analyses, taking concrete steps toward the development of an explainable run-time monitor. More specifically, conformance checking algorithms can pinpoint local diagnoses when comparing run-time control-flow with the prescriptive process model built with process discovery, supporting the root cause analysis of abnormal behavior.
Process mining has been applied for anomaly detection in several domains, including business processes of organizations [bezerra2013adlogspais, vitale2025cfad], software applications [cinque2019hiddenapplicationerrorspm, pecchia2020applicationfailuresanalysispm], computer networks [saintpierre2014dnspmad, hemmer2021adpm], and cyber-physical systems [myers2018icsadpm, vitale2025pmdt]. Notably, process mining has been also applied using an ERTMS/ETCS dataset generated from a high-level description of one use-case scenario in [debenedictis2023dtadiiot, vitale2025cfad]. However, previous work did not thoroughly analyze ERTMS/ETCS L2 use cases or investigate process mining support for their development, monitoring, and the explanation of control-flow deviations from prescriptive specifications. In contrast, we propose a dedicated methodology that systematically combines behavior characterization, run-time monitoring, and anomaly detection and explanation. Our approach leverages offline ERTMS/ETCS L2 simulation or execution to extract a high-fidelity execution model of the target use case, integrates run-time execution with online conformance checking, applies machine learning-based post-processing of conformance diagnoses using clustering algorithms, and ultimately labels and localizes anomalies to pinpoint the faulty ERTMS/ETCS L2 components.
@@ 2.1 Preliminary Definitions
Traditional definitions of dependability typically emphasize either “the ability to deliver service that can justifiably be trusted” or “the ability to avoid service failures that are more frequent and more severe than is acceptable” [avizienis2004dependabilitytaxonomy]. These definitions encompass a broad range of strategies for achieving dependability, including fault prevention, fault tolerance, fault removal, and fault forecasting. However, they do not explicitly address a key concept that is central to this paper: resilience. Resilience refers to a system’s ability to continue delivering correct service in the face of changes in its run-time environment, including functional, environmental, and technological changes [laprie2005resilience].
To achieve resilience in modern computer-based railways, including those based on ERTMS/ETCS L2, we propose to add specific software monitors performing run-time control-flow anomaly detection. Anomaly detection is a broad research area that aims to find “patterns in data that do not conform to a well defined notion of normal behavior" [chandola2009ad]. Different types of anomaly detection address different types of anomalies. In this paper, we address control-flow anomalies, namely those anomalies that manifest as unknown, skipped, and wrongly-ordered activities of event logs from computer-based systems [vitale2025cfad]. In addition, we aim to perform control-flow anomaly detection at run-time, including when the system is operating in non perfectly known/predictable conditions, hence subject to functional, environmental, and technological uncertainties. Finally, it is worth mentioning that we propose an unsupervised approach to anomaly detection and localization, as we aim to characterize normal behavior from reference execution traces of the target system, and localize anomalies based on the model-based diagnoses we perform against a reference process model of the software.
@@ 2.2 ERTMS/ETCS L2
ERTMS/ETCS is characterized by a certain level of operation, from 1 to 3, which defines the degree of automation performance, and the required trackside and on-board equipment [ghazel2014formalizingertmsetcsspecs].
@@
In this paper, we focus on the most successful level of operation to date, which is level 2. At this level, the on-board system collects the train’s position through Eurobalises, which are electronic beacons installed across the tracks. The PR is forwarded to the RBC, which is responsible for train separation; the RBC elaborates the MA based on track occupation. Fig. 1 shows a common use-case in ERTMS/ETCS, namely the MA request and provision to supervise train traffic. This use-case can be implemented within the Start-of-Mission procedure, which is one of the procedures documented in specification SUBSET-026; such a subset prescribes the behavior of ERTMS/ETCS constituents in all reference operational scenarios. Each step, labeled 1 to 7, requires either exchanging information between the on-board equipment of the requesting and preceding trains with the RBC, or using the on-board modules to perform some kind of data processing. The most important task performed on-board is the calculation of the so-called braking curve (also known as dynamic speed profile).
We assume that several types of run-time anomalies may occur across the many modules that compose ERTMS/ETCS L2, such as unexpected train data exchanged between the Driver Machine Interface (DMI) and the RBC through the Radio Transmission Module (RTM), mispositioning of balises, and issues in establishing connections with the RBC [ditommaso2005anomaliessimulation, valdivia2017etcssafetytesting, cai2019reliabilityavailabilityeval]. These anomalies may propagate across different components, and may influence other key modules, such as the European Vital Computer (EVC), which is the on-board component responsible for performing critical operations, such as calculating the braking curve.
@@ 2.3 Process Mining and Machine Learning
Process mining “aims to improve operational processes through the systematic use of event data” by mainly two types of algorithms: process discovery and conformance checking [aalst2022pmhb]. Process discovery attempts to build a normative behavioral model from historical event data collected by monitoring a reference application or process. Conformance checking aligns new event data to the normative model to diagnose any deviations from the control flow captured by the model. In addition, conformance checking can be performed online, i.e., it can be applied to streaming event data through, e.g., window-based approaches and prefix alignments [burattin2020streamingpdcc, vanzelst2019onlinecc].
Machine learning allows encoding data patterns into models. Traditional machine learning approaches capture patterns in data through rule-based learning, probabilistic modeling, linear and non-linear modeling, and clustering [witten2016mltools]. Advanced techniques include deep learning, which implements (deep) artificial neural networks mimicking the functioning of the human brain [dong2021dlsurvey]. Machine learning algorithms also differ based on their learning paradigm (supervised, unsupervised, and semi-supervised) and degree of explainability [rawal2022trustworthyai].
Previous works have shown that conformance checking can be combined with unsupervised machine learning for explainable control-flow anomaly detection. The approach proposed in this paper is inspired by the work reported in reference [vitale2025cfad], and proposes performing unsupervised run-time monitoring of ERTMS/ETCS L2 control flow through the combination of conformance checking and clustering.
@@ 3.1.1 Offline ERTMS/ETCS Execution
The first step begins with offline ERTMS/ETCS execution, where a test suite is applied to a simulation environment or an actual system implementation. The simulation of ERTMS/ETCS systems can provide different types of high-level events that can be used to characterize the execution model of the target scenario. For example, di Tommaso et al. [ditommaso2005anomaliessimulation] simulated the Track Ahead Free (TAF) procedure with a trackside simulation environment. In this procedure, the RBC sends a TAF request to the train driver through the DMI to possibly grant a full supervision MA. Gaspari et al. [gaspari2019hybrideuropeanrailtrafficmgms] exemplified another remarkable simulation approach, involving the execution of abstract state machines representing several ERTMS/ETCS L3 scenarios to test the correctness of their models. The execution traces are collected to model the software behavior. We define a set of $k$ traces as $\Sigma=\{\sigma_{1},\sigma_{2},\dots,\sigma_{k}\}$, where each trace $\sigma$ is an ordered sequence of events: $\sigma=\langle e_{1},e_{2},\dots,e_{|\sigma|}\rangle\in\Sigma$, with $e_{j}$ representing the $j$-th event. The definition of an event depends on the level of abstraction used in simulating ERTMS/ETCS L2. Since these systems involve distributed software components communicating over a railway network, events may correspond to specific software procedures executed by these components. This is a realistic representation, as safety-critical ETCS software is usually implemented in C/C++ [angeletti2010safetycriticalsoftwarecoverageanalysis, cimatti2012ertmsl2verificationvalidation, karg2016modeldrivenseopenetcs].
@@ 3.1.2 Model Building
After the collection of execution traces, the model building activity applies a process discovery algorithm $\gamma$ to find an execution model $M$ from the set of traces representing the actual behavior of the system, i.e., $\gamma(\Sigma)=M$. There are many process discovery algorithms classified by their different learning approaches, including footprint-based and directly-follows-graph-based approaches that uncover local relationships between events (e.g., the $\alpha$-miner and split miner), divide-and-conquer approaches that iteratively find relationships starting from the most general one (e.g., the inductive miner and partially ordered workflow language miner), and region-based approaches that obtain so-called regions from either state-based or language-based represesentations of the input log (e.g., the integer linear programming-based miner) [aalst2022pmhb]. Each process discovery algorithm has its own representational bias, which influences the quality of the resulting execution model. The execution model can be, e.g., a Petri net, a Business Process Modeling Notation (BPMN) model, or a process tree. The target formalism depends on the requirements of the applications. For example, a Petri net may be better suited for safety-critical environments due to its formal semantics. Hence, we consider the Petri net as the target formalism in the following.
@@ 3.2.1 Online ERTMS/ETCS Execution
The second step is initiated by the online execution of an ERTMS/ETCS procedure with an unknown input. The real process is monitored throughout its execution. Such run-time monitoring can be performed in different ways. A plausible mechanism would be source-code instrumentation by well-defined logging rules, as proposed in [cinque2013rulebasedmonitoring]. This technique places logging rules to record specific events occurring in a program, such as interactions between objects, procedure calls, and the invocation of microservices. Its feasibility is due to the availability of source code in ERTMS/ETCS L2 implementations, as these systems require thorough white-box analyses to certify their safety levels. Fig. 3 shows an example deployment diagram where the on-board and trackside subsystems communicate with each other and send run-time events to a monitor server. These run-time events configure a partial trace, and are checked against the execution model by a fault detector using a conformance checker that implements process mining techniques.